7.5 mm 9-76 The ladder is supported on the rough surface at A and by a smooth wall at B. If a man weighing 675 N stands upright at C. determine the principal stresses in one of the legs at point D. Each leg is made from a 25 mm-thick board having a rectangular cross section. Assume that the total weight of the man is exerted vertically on the rung at Cand is shared equally by each of the ladder's two legs Neglect the weight of the ladder and the forces developed by the man's arms m75 mmm 12 m
d Nasi 4S48 9-2 The state of stress at a point in a mcnber is shown n the clement. Determine the stress components acting the inclined plane AB. Solve the problem using the method of equilihriom descried in Sec. 9.1 5 MP 3 MP .3. The state of stress at ap ut in a member n shown on the clement. Determin the stres componcats acting on the inclined plane AB.Solve the problem using the me 9-44 Solve Prob. 9-3 using Mohr's circke 9-6 Detcrmine the normal stress and shear stress acting on the inclined planc AB. Solve the problem using the method of equildheius described in Sec.9.1 MP 9-45. Solve Prob. 9-6 using Mohr's circle. 9-27. The bracket is subjccted to the fouve of 15kN Detcrmine the principal stress and maximum meplane shear stress at pont Bon the cross sectional section ISIN specily the aricntativn of ths state of stres rsalts on clements o ma
Answer #1

-21-63 오 15 12. 21-63 ho 2 4 v = 8.89 21.6.3軀 6198 MN 2-2 5 て = va - 8.88 (t) 3-34-1 Ro.ST 6 C (So 285, 6 Co, 3-34-7) -124.57SMPg てxyı = ΔΑ